1000 tickets were sold. Adult tickets cost $8.50,children's cost $4.50,and a total of $7300 was collected. How many of each kind were sold?
Answer by richwmiller(17219)

x+y=1000,
8.5*x+4.5*y=7300
x=1000-y
8.5*(1000-y)+4.5*y=7300
8500.0-8.5y+4.5*y=7300
-4.0*y=-1200.0
y=300.0
x=1000-y
x=700 at 8.50 and y=300 at 4.50
check
700+300=1000
8.5*x+4.5*y=7300
8.5*700+4.5*300=7300
5950.0+1350.0=7300
7300.0=7300
ok
codetick